# Canonical Policy: Learning Canonical 3D Representation for Equivariant Policy

### v1

@@ IV Method
In this section, we first present the theoretical foundation of canonical representations for 3D observations. We then introduce canonical policy framework, which enables SE(3)-equivariant visual imitation by mapping observations, robot states, and actions into a shared canonical frame. Finally, we describe the point cloud encoder designed for feature extraction.
IV-A Theory of Canonical Representations
Assumption 1.
A set of 3D point clouds can be partitioned into a finite number of mutually exclusive subsets, denoted as, $\mathcal{D}=\{\mathcal{G}_{1},\mathcal{G}_{2},\cdots,\mathcal{G}_{p}\},$ such that
$\forall\mathbf{R}\in\mathbb{R}^{3\times 3},\ \forall\mathbf{b}\in\mathbb{R}^{3\times 1},\ \forall i\neq j,\quad\mathcal{G}_{i}\cap(\mathbf{R}\mathcal{G}_{j}+\mathbf{b})=\emptyset,$ (1)
where $p$ is the total number of groups, and $\mathbf{R}$ and $\mathbf{b}$ denote rotation matrix and translation vector, respectively. Moreover, within each group, we have
Remark 1.
Under Assumption 1, we can see that $\mathcal{G}_{i},i=1\ldots p$ forms an equivariant group, such that for any rigid-equivariant function $f$, the following holds:
In the context of visual-conditioned policy learning, it is reasonable to assume that new observations from the same task can be related to samples in $\mathcal{D}$ through rigid transformations, which is summarized below.
Assumption 2.
For a set of point clouds $\mathcal{D}$ collected from a demonstrated task, any point cloud observation from the same task but outside $\mathcal{D}$ can be mapped to a point cloud within $\mathcal{D}$ through a rigid transformation.
In practice, the collected dataset $\mathcal{D}$ is typically diverse enough to cover a wide range of variations. Therefore, unseen point cloud observation but from the same task can be approximately mapped to $\mathcal{D}$ by rigid transformations. For example, in robotic manipulation tasks, a dataset may include multiple demonstrations of grasping the same object from different orientations and viewpoints. These demonstrations inherently capture variations such as rotations, translations, and perspective shifts, all of which can be effectively modeled as rigid body transformations in 3D space. Consequently, unseen geometric layouts can often be mapped to known examples in the dataset through appropriate rigid transformations.
To simulate observations with new poses not present in the original dataset, we define a transformed variant using rigid transformations. Specifically, let $\mathbf{R}_{i}\in\mathbb{R}^{3\times 3}$ and $\mathbf{b}_{i}\in\mathbb{R}^{3\times 1}$ be randomly sampled rotation matrices and translation vectors. We define: $\mathcal{D}^{\mathrm{rd}}=\{\mathcal{G}_{1}^{\mathrm{rd}},\mathcal{G}_{2}^{\mathrm{rd}},\cdots,\mathcal{G}_{p}^{\mathrm{rd}}\}$, where $\mathcal{G}_{i}^{\mathrm{rd}}=\mathbf{R}_{i}\mathcal{G}_{i}+\mathbf{b}_{i}.$ Here, the superscript “rd” denotes randomly transformed representation.
Notably, Assumptions 1 and 2 still hold for $\mathcal{D}^{\mathrm{rd}}$. Moreover, in most cases, data points generated by randomly sampled transformations are considered out-of-distribution with respect to the original dataset. Based on this observation, we propose leveraging the equivariance property to map both in-distribution and out-of-distribution samples to a canonical distribution, which serves as a standardized representation for each group: $\mathcal{D}^{\mathrm{cn}}=\{{\mathcal{G}}_{1}^{\mathrm{cn}},{\mathcal{G}}_{2}^{\mathrm{cn}},\cdots,{\mathcal{G}}_{p}^{\mathrm{cn}}\},$ where the superscript “cn” denotes canonical representation. Each ${\mathcal{G}}_{i}^{\mathrm{cn}}$ denotes the canonical representation shared by both $\mathcal{G}_{i}$ and its transformed counterpart $\mathcal{G}_{i}^{\mathrm{rd}}$, such that
$\forall\mathbf{x}\in\mathcal{G}_{i}\cup\mathcal{G}_{i}^{\mathrm{rd}},\ \exists{\mathbf{x}}^{\mathrm{cn}}\in{\mathcal{G}}_{i}^{\mathrm{cn}},\ \exists\mathbf{R},\mathbf{b},\quad\mathbf{x}=\mathbf{R}{\mathbf{x}}^{\mathrm{cn}}+\mathbf{b}.$
Since Assumption 1 holds for both $\mathcal{G}_{i}$, $\mathcal{G}_{i}^{\mathrm{rd}}$, and ${\mathcal{G}}_{i}^{\mathrm{cn}}$ and Remark 1 implies that they belong to the same equivariant group, for any equivariant function $f$, the following holds:
$\forall\,\mathbf{x}\in\mathcal{G}_{i}\cup\mathcal{G}_{i}^{\mathrm{rd}},\ \exists\,\mathbf{x}^{\mathrm{cn}}\in\mathcal{G}_{i}^{\mathrm{cn}},\ \exists\,\mathbf{R},\mathbf{b},$
$f(\mathbf{x})=\mathbf{R}f(\mathbf{x}^{\mathrm{cn}})+\mathbf{b}.$ (3)
This implies that the equivariance property of $f$ allows rigid transformations to be factored out from the original dataset, which can in turn benefit the training of a robot policy $\pi$. By leveraging this property and estimating the transformation parameters $\mathbf{R}$ and $\mathbf{b}$ as described in Equation (3), the policy $\pi$ can generalize any observation within the equivariant group as
$\forall\mathbf{x}\in\mathcal{G}_{i}\cup\mathcal{G}_{i}^{\mathrm{rd}},\ \exists{\mathbf{x}}^{\mathrm{cn}}\in{\mathcal{G}}_{i}^{\mathrm{cn}},\ \exists\mathbf{R},\mathbf{b},$
$\pi(\mathbf{x})=\mathbf{R}\pi({\mathbf{x}}^{\mathrm{cn}})+\mathbf{b}.$
Consequently, both in-distribution and out-of-distribution observations, whether from $\mathcal{D}$ or $\mathcal{D}^{\mathrm{rd}}$, can be reliably mapped to a canonical form. The policy can thus operate consistently on the canonical representation. The only requirement is to account for the rigid transformation between the original and canonical representations. Once this transformation is resolved, all inputs including out-of-distribution observations effectively become in-distribution inputs, allowing the policy to achieve generalizable performance.
IV-B Differences of Canonical Representation and Data Augmentation
Figure 1: Comparison of canonical representation and data augmentation: (a) data augmentation; (b) data canonicalization
To better understand the benefits of canonical representations, we provide a conceptual comparison with data augmentation, which is a commonly used technique for improving generalization to out-of-distribution data.
As illustrated in Fig. 1(a), data augmentation operates by applying random rigid transformations to the input data. These augmented samples are then used to extract the observation features, with the hope that exposure to diverse poses can improve robustness at inference time. In contrast, Fig. 1(b) shows that canonical representation explicitly normalizes inputs to a consistent pose, regardless of their original transformation, before passing them to the downstream network.
However, data augmentation has fundamental limitations in pose generalization. Since the augmented samples are finite and sampled from a subset of all possible rigid transformations, the trained policy may fail to generalize to unseen poses that were not covered during training. This leads to degraded performance when encountering new viewpoints or spatial configurations of the same object. In contrast, the canonical representation approach addresses this limitation by utilizing an equivariant mapping that transforms any instance in the same equivariant group into a unique canonical pose. As a result, the network always observes the same canonicalized input, enabling better generalization even to unseen but transformation-equivalent inputs.
Formally, canonical representation can be seen as a form of pre-alignment that enforces invariance to a certain group of transformations. Instead of requiring the network to learn this invariance implicitly through data diversity, the equivariant transformation module explicitly reduces the input variation. This enable the policy to learn action mappings from a standardized input space.
IV-C Canonical Policy Pipeline
This section outlines the architecture of canonical policy, which achieves transformation-invariant action prediction by operating in a canonical coordinate frame. The core idea is to map both visual and low-dimensional inputs (e.g., robot state and action) to a consistent pose, reducing geometric variation before policy inference. This allows the policy to focus on task-relevant behavior rather than handling pose differences. We elaborate on each component of this framework in the following subsections.
As shown in Fig. 2, given a point cloud observation $\mathbf{x}\in\mathcal{G}_{i}$, we first compute its mean $\mathbf{x}^{\mathrm{mn}}\in\mathbb{R}^{3\times 1}$ and subtract it from each point to remove the effect of translation, where the superscript “mn” indicates the mean operator. This centering step aligns the point cloud at the origin. Next, an SO(3)-equivariant network is used to compute the relative rotation $\mathbf{R}\in\mathrm{SO}(3)$ with respect to its canonical pose. The details of this process will be discussed in the next subsection.
After obtaining the canonical pose for each input point cloud, the canonical form is fed into the Point Cloud Aggregation Encoder Block (see the bottom of Fig. 2). Within this block, local neighborhood features are extracted and processed through multiple fully connected layers to generate features for the canonicalized point cloud.
For robot state inputs and actions, we use end-effector positions and orientations, which are widely adopted in visual imitation learning [1, 9, 3]. In this setting, a data canonicalization pipeline analogous to that used for point clouds can be applied, aligning both the state and action to a canonical frame via the estimated $\mathrm{SE(3)}$ transformation. It is worth noting that our method operates in Cartesian space rather than joint space representations (e.g., joint angles [19]).
In practice, the orientation component of both the robot state and action is not represented as a rotation matrix. The robot state typically encodes orientation using unit quaternions, which we convert to rotation matrices before applying the canonical transformation. Similarly, the format of action orientation depends on the control mode: in absolute control, it is represented using the 6D continuous rotation formulation [44], while in relative control, it is specified as an axis-angle vector. Both formats are first converted to rotation matrices before applying the transformation and converted back afterward for prediction. For simplicity, we omit format conversions in the subsequent equations and use rotation matrices to represent $\mathrm{SO}(3)$ rotations. Additionally, to streamline the derivation, we consider only a single robot state from the observation sequence $\mathcal{O}$ of length $m$, and a single action from the action sequence of length $n$ in the subsequent discussion.
$\mathbf{S}_{t}^{\mathrm{cn}}=\mathbf{T}_{t}^{-1}\mathbf{S}_{t}.$
Absolute Control
${\mathbf{A}}_{k,t}^{\mathrm{cn}}=\mathbf{T}_{t}^{-1}\mathbf{A}_{k,t}.$ (4)
$\mathbf{A}_{k-1,t}=\mathbf{T}_{t}{\mathbf{A}}_{k-1,t}^{\mathrm{cn}}.$
Relative Control
$\mathbf{A}_{k,t}=\Delta\mathbf{A}_{k,t}\cdot\mathbf{A}_{k,t-1}.$ (5)
$\mathbf{A}_{k,t}^{\mathrm{cn}}$ $=\mathbf{T}_{t}^{-1}\mathbf{A}_{k,t}$
$=\mathbf{T}_{t}^{-1}\Delta\mathbf{A}_{k,t}\cdot\mathbf{A}_{k,t-1}$
Since consecutive observations at high sampling rates, such as 10 Hz [14, 3], typically exhibit only minor differences, particularly in smooth and continuous manipulation behaviors [1], we approximate the canonical transformations at adjacent timesteps as equal, i.e., $\mathbf{T}_{t-1}\approx\mathbf{T}_{t}$. Accordingly, Equation (6) becomes:
Comparing this result with the canonical relative action update:
IV-D SE(3) Equivariance via Data Canonicalization
@@
The formulation in Section IV-C builds upon the availability of a canonical pose for each observation, enabling unified representation of all data for downstream policy learning. In this section, we formally derive how such a canonical pose can be computed via an $\mathrm{SE}(3)$-equivariant data canonicalization procedure. Specifically, we detail how a rotation matrix is derived using an SO(3)-equivariant network, which allows point clouds related by rigid transformations to be mapped to a consistent canonical frame.
Figure 3: Data canonicalization pipeline. Input point clouds are centered and processed by a shared $\mathrm{SO}(3)$-equivariant encoder to estimate rotation matrices. The inverse of the estimated rotation is then applied to align each observation to a shared canonical representation. Superscripts “de”, “mn”, and “cn” denote the decentered representation, the mean operator, and the canonical representation, respectively.
@@
where the superscript “de” denotes the decentered representation. This transformation effectively eliminates the effect of translation, making it translation invariance, ensuring that the relationship between any two decentered elements is solely determined by a rotation.
@@
A natural instantiation of such an SO(3)-equivariant network is the Vector Neuron framework [45], as it operates directly on vector features rather than scalars. Specifically, Vector Neuron represents each feature as a 3D vector and aggregates them into a $3\times D$ tensor. Although each column corresponds to a distinct equivariant feature component, they are not transformed independently. Instead, the entire tensor undergoes a shared transformation under a single rotation matrix, preserving global rotational equivariance. As a result, Equations (7) and (8) hold by design.
@@
This procedure, illustrated in Fig. 3, ensures that no matter which equivariant group the observation point cloud belongs to: 1) The input is first decentered by subtracting its mean, ensuring translation invariance. 2) It is then passed through an SO(3)-equivariant network to estimate a corresponding rotation matrix. 3) Finally, the input is derotated using this matrix to produce a unique canonical pose representative of its equivariant group.
In this case, the data within the same group will eventually convert to the identical representation. If the policy is trained on such a canonical pose, then when it encounters a completely new data that is a rotated or translated version of a previously seen sample, the canonical mapping step will transform it into the same canonical form of the seen one. Therefore, the policy can seamlessly generalize across variations in rotation and translation, significantly improving its robustness.
Furthermore, after computing the mean vector and estimated rotation matrix for each sample, we transform the robot state and noisy action using the SE(3) inverse and forward procedures described above. This guarantees alignment between the low-dimensional data and the transformed observed point cloud.
@@
[…]
@@
IV-E Point Cloud Aggregation Encoder
@@
Building on the canonical observations described in the previous section, we introduce a point cloud aggregation encoder designed to extract informative features by leveraging the spatial structure inherent to point cloud inputs in the canonical frame. This facilitates robust policy learning from geometrically consistent inputs.
@@
DP3 [3] explores various point cloud encoders that have demonstrated effectiveness in computer vision tasks, including PointNet++ [46], PointNeXt [47], and Point Transformer [48]. Their findings suggest that a simple multi-layer perceptron (MLP) is a highly effective approach for point cloud encoding in imitation learning. In addition to MLPs, we also examine alternative encoders such as DGCNN [49], with a detailed comparison presented in the following section. Based on these insights, we adopt a simple MLP architecture with residual connections, facilitating the training of deeper networks for robust feature extraction.
@@
To exploit this property, we first identify the neighbors of each point in the point cloud before applying MLP-based feature extraction. PointMLP [50] introduces a geometric affine module that effectively aggregates information from neighboring points, as illustrated in Fig. 4.
@@
[…]

### v2

@@ IV Method
In this section, we first present the data canonicalization procedure, which estimates the $\mathrm{SE}(3)$ transformation that maps each sample to a canonical frame. We then introduce the canonical policy framework, which achieves $\mathrm{SE}(3)$-equivariant visual imitation by mapping observations, robot states, and actions into a shared canonical frame. Finally, we describe the point cloud encoder designed for feature extraction.
IV-A SE(3) Equivariance via Data Canonicalization
@@
The formulation in Section III-C builds the theoretical foundation of canonical representations for 3D observations, enabling the policy to generalize to unseen observations that are SE(3)-equivariant to those seen during training. The key to canonical representation is determining the transformation parameters in Eq. (4). Therefore, in this section, we formally derive how such a canonical pose can be computed via an $\mathrm{SE}(3)$-equivariant canonicalization procedure. Specifically, we describe how the rotation matrix is estimated using an $\mathrm{SO}(3)$-equivariant network, enabling point clouds related by rigid motions to be mapped to a consistent canonical frame.
@@
where the superscripts “de” and “mn” denote the decentered representation and the mean operator, respectively. This transformation effectively eliminates the effect of translation, making it translation invariance, ensuring that the relationship between any two decentered elements is solely determined by a rotation.
@@
A natural instantiation of such an SO(3)-equivariant network is the Vector Neuron framework [VN], as it operates directly on vector features rather than scalars. Specifically, Vector Neuron represents each feature as a 3D vector and aggregates them into a $3\times D$ tensor. Although each column corresponds to a distinct equivariant feature component, they are not transformed independently. Instead, the entire tensor undergoes a shared transformation under a single rotation matrix, preserving global rotational equivariance. As a result, Equations (5) and (6) hold by design.
@@
Figure 3: Vector Neuron framework for estimating $\mathrm{SO}(3)$-equivariant rotation matrix. A graph is built from the decentered point cloud, local features are aggregated into global equivariant features, and Schmidt orthogonalization generates a rotation in $\mathrm{SO}(3)$ or $\mathrm{SO}(2)$ that aligns the input to a canonical frame.
In this case, the data within the same group will eventually convert to the identical representation. If the policy is trained on such a canonical pose, then when it encounters a completely unseen data that is a rotated or translated version of a previously seen sample, the canonical mapping step will transform it into the same canonical form of the seen one. Therefore, the policy can seamlessly generalize across variations in rotation and translation, significantly improving its robustness.
@@
We extract the in-plane rotation angles by projecting $\mathbf{R}_{\mathbf{x}}$ and $\mathbf{R}_{\mathbf{y}}$ onto the $xy$-plane and applying the $\mathrm{atan2}$ function:
$\theta_{\mathbf{x}}$ $=\mathrm{atan2}(y_{1},x_{1}),$
$\theta_{\mathbf{y}}$ $=\mathrm{atan2}(x_{1}\sin\theta+y_{1}\cos\theta,\ x_{1}\cos\theta-y_{1}\sin\theta).$
Using the rotational property of $\mathrm{atan2}$, we observe: $\theta_{\mathbf{y}}=\theta_{\mathbf{x}}+\theta.$ Then, the canonicalized form of $\mathbf{y}^{\mathrm{de}}$ becomes:
@@
Figure 3 illustrates the detailed pipeline for estimating the $\mathrm{SO}(3)$-equivariant rotation matrix using the Vector Neuron architecture. The network takes the decentered point cloud $\mathbf{x}^{\mathrm{de}}$ as input and then constructs a k-nearest neighbor (KNN) graph to capture local geometric structure and derive features such as relative offsets, absolute positions, and local directions, thereby enhancing robustness to noise. Stacked $\mathrm{SO}(3)$-equivariant linear layers and activation blocks with mean pooling yield a global equivariant feature $\mathbf{F}_{\text{equiv}}$, which is then used to compute two $\mathrm{SO}(3)$-equivariant vectors $\{\mathbf{r}^{1}_{\mathbf{x}},\,\mathbf{r}^{2}_{\mathbf{x}}\}$. Applying Schmidt orthogonalization to these vectors produces an $\mathrm{SO}(3)$ rotation matrix. For tasks with only an $\mathrm{SO}(2)$ ambiguity, the network further extracts the in-plane (z-axis) rotation. This rotation maps $\mathbf{x}^{\mathrm{de}}$ to its canonicalized counterpart.
The reliability of the $\mathrm{SO(3)}$-equivariant network under noisy conditions, parameter sensitivity, and feature visualization will be systematically examined in Section V.
IV-B Canonical Policy Pipeline
Following the data canonicalization described in Section IV-A, this section presents the architecture of the canonical policy, which performs transformation-invariant action prediction by operating in a canonical coordinate frame. The core idea is to map both visual and low-dimensional inputs (e.g., robot proprioceptive state and action) to a consistent pose, reducing geometric variation before policy inference. This allows the policy to focus on task-relevant behavior rather than handling pose differences.
As shown in Fig. 4, given a point-cloud observation $\mathbf{x}\in\mathcal{G}_{i}$ for which the canonical pose $\mathbf{x}^{\mathrm{cn}}$ and the corresponding transformation have been obtained via the data canonicalization procedure, the canonical form is fed into the Point Cloud Aggregation Encoder Block (see the bottom of Fig. 4). Within this block, local-neighborhood features are extracted and processed through multiple fully connected layers to produce features for the canonicalized point cloud. The encoder design and hyperparameters are described in the next subsection.
For robot state inputs and actions, we use end-effector positions and orientations, which are widely adopted in visual imitation learning [diffusion_policy, equidiff, DP3]. In this setting, a data canonicalization pipeline analogous to that used for point clouds can be applied, aligning both the state and action to a canonical frame via the estimated $\mathrm{SE(3)}$ transformation. It is worth noting that our method operates in Cartesian space rather than joint space representations (e.g., joint angles [zhao2024aloha]).
In practice, the orientation component of both the robot state and action is not represented as a rotation matrix. The robot state typically encodes orientation using unit quaternions, which we convert to rotation matrices before applying the canonical transformation. Similarly, the format of action orientation depends on the control mode: in absolute control, it is represented using the 6D continuous rotation formulation [zhou2019continuity], while in relative control, it is specified as an axis-angle vector. Both formats are first converted to rotation matrices before applying the transformation and converted back afterward for prediction. For simplicity, we omit format conversions in the subsequent equations and use rotation matrices to represent $\mathrm{SO}(3)$ rotations. Additionally, to streamline the derivation, we consider only a single robot state from the observation sequence $\mathcal{O}$ of length $m$, and a single action from the action sequence of length $n$ in the subsequent discussion.
$\mathbf{S}_{t}^{\mathrm{cn}}=\mathbf{T}_{t}^{-1}\mathbf{S}_{t}.$
Absolute Control
${\mathbf{A}}_{k,t}^{\mathrm{cn}}=\mathbf{T}_{t}^{-1}\mathbf{A}_{k,t}.$ (10)
$\mathbf{A}_{k-1,t}=\mathbf{T}_{t}{\mathbf{A}}_{k-1,t}^{\mathrm{cn}}.$
Relative Control
$\mathbf{A}_{k,t}=\Delta\mathbf{A}_{k,t}\cdot\mathbf{A}_{k,t-1}.$ (11)
$\mathbf{A}_{k,t}^{\mathrm{cn}}$ $=\mathbf{T}_{t}^{-1}\mathbf{A}_{k,t}$
$=\mathbf{T}_{t}^{-1}\Delta\mathbf{A}_{k,t}\cdot\mathbf{A}_{k,t-1}$
Since consecutive observations at high sampling rates, such as 10 Hz [wang2024dexcap, DP3], typically exhibit only minor differences, particularly in smooth and continuous manipulation behaviors [diffusion_policy], we approximate the canonical transformations at adjacent timesteps as equal, i.e., $\mathbf{T}_{t-1}\approx\mathbf{T}_{t}$. Accordingly, Equation (12) becomes:
Comparing this result with the canonical relative action update:
IV-C Point Cloud Aggregation Encoder
@@
Having established the canonicalized observations in the previous subsection, we now introduce a point cloud aggregation encoder designed to extract informative features by leveraging the spatial structure inherent to point cloud inputs in the canonical frame. This facilitates robust policy learning from geometrically consistent inputs.
@@
DP3 [DP3] explores various point cloud encoders that have demonstrated effectiveness in computer vision tasks, including PointNet++ [pointnet++], PointNeXt [pointnext], and Point Transformer [pointtransformer]. Their findings suggest that a simple multi-layer perceptron (MLP) is a highly effective approach for point cloud encoding in imitation learning. In addition to MLPs, we also examine alternative encoders such as DGCNN [DGCNN], with a detailed comparison presented in the following section. Based on these insights, we adopt a simple MLP architecture with residual connections, facilitating the training of deeper networks for robust feature extraction.
@@
To exploit this property, we first identify the neighbors of each point in the point cloud before applying MLP-based feature extraction. PointMLP [PointMLP] introduces a geometric affine module that effectively aggregates information from neighboring points, as illustrated in Fig. 5.
@@
First, the neighborhood of each point is determined based on its $\mathit{xyz}$ coordinates, yielding a set of neighboring features represented as $\mathcal{F}_{i}=\{\mathbf{F}_{i}^{1},\mathbf{F}_{i}^{2},\dots,\mathbf{F}_{i}^{q}\}$, where $i$ denotes the $i$th point and $q$ represents the number of neighbors. Within this local neighborhood, we compute the mean $\mu$ and standard deviation $\sigma$ of the features and apply feature normalization. To further enhance robustness in feature aggregation, affine transformation parameters $\alpha$ and $\beta$ are applied to the normalized local features, producing the final aggregated features $\mathcal{F}_{i}^{\mathrm{ag}}$: